White dot for spacing only
The Dice Project


Operational Meeting

Minutes of the meeting held on Wednesday 8th February 2017 in IF-4.31

Present:
Ross Armstrong, Richard Bell, Toby Blake, Lindsey Brown, Neil Brown, Roger Burroughes, Chris Cooke, Alison Downie, Ian Durkacz, Graham Dutton, Jennifer Oxley, Iain Rae, George Ross (convener and minutes), Alastair Scobie, Craig Strachan
  1. Minutes of the last meeting

    These were accepted as a reasonable account of the last meeting.

  2. Actions and blog articles

    Actions discussed:

    Actions added or revived:

    Actions deferred:

    Actions completed:

    Blog articles suggested:

  3. Report from Computing Executive Group

    1. Appleton Tower
    2. CCPAG verbal report
      • Network strategy review. Tony and Sam reported. Alastair is planning a presentation at a Development meeting.
      • Lynda.com
      • SciCOs are looking for talks
      • The JCMB CSR OLA is almost done!
      • The Windows 10 desktop hasn't had enough effort available. Angus and Myles have been seconded to it.
      • Lecture capture (see also the US Unit report): the project is on time. There is some ongoing discussion as to exactly what will be captured and what won't. (For example, some lecturers make heavy use of blackboards.)
      • KB power outages: a business continuity plan is going to be formulated, which is expected to cover at least how information is communicated during outages.
      • There's going to be a "tidy" of functional accounts. What do we need? What could be deleted? No-one in the room had heard anything about this proposal, though apparently the target is March some time. There was some discussion as to how we (and presumably others) obtain new functional accounts or spot and use existing useful-looking ones, and how this might interact with IS's process. Toby will coordinate our repsponse.
      • The remit of the EdWeb governance group is apparently going to be extended to cover all web servers throughout the University. There was some doubt as to how this would work in practice.
      • The research only OpenStack service had been announced. It doesn't appear to allow us to try it out so that we can better inform our users.
      • The new server-specification round has started. Iain volunteered to be our representative this time.
      • Data-centre strategy: an oversight group is being created, to include EPCC and us.
      • Our loan UPS for the CSR PoP switch stack was agreed. Buying a new one for the purpose wasn't discussed.
      • (Recent EdWeb degradation wasn't discussed. It appears to have been a load-balacer capacity issue, rather than a problem with the service per se.)
  4. Reports from Units

  5. Topics for discussion

  6. Items for the Computing Systems blog

    Domain squatters (George).

  7. AOCB

  8. Next meeting

    The next meeting will be on Wednesday 22nd February at 10:00, location to be confirmed, chaired by George.


 : Operational : Meetings 

Mini Informatics Logo - Link to Main Informatics Page
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh
Spacing Line